# HodgeFormer: Transformers for Learnable Operators on Triangular Meshes through Data-Driven Hodge Matrices

### v1

@@ Abstract
Index Terms:
3D shape analysis, learnable operators, representation learning
I Introduction
@@
In this work, we introduce a novel Transformer architecture inspired by the explicit construction of the Laplacian operator in Discrete Exterior Calculus (DEC) adjusted to triangular 3-dimensional meshes. This kind of connection is not novel by itself and was the main idea of HodgeNet by [5], where input features were used for learning Hodge Star operators as diagonal matrices. These matrices were then used for constructing Laplacian operators, which were then applied to input features in a spectral learning kind of approach. Specifically, we draw connections between discrete Hodge Star operators and Transformer-style attention mechanisms. While most common constructions of discrete Hodge Star operators are limited to diagonal matrices, these represent only some specific Hodge operator realizations. In fact, the Galerkin method, widely used in Finite Element Methods (FEM) literature [6], prescribes systematic ways to discretize various differential operators, including the Hodge operator. The key idea is to project onto carefully designed basis functions (e.g., piecewise linear or higher order elements) and account for the overlap of these basis functions. These overlaps are used to design the mass matrix, which acts as a discrete Hodge operator. Unlike the diagonal Hodge operator, which considers only self-contributions, the Galerkin Hodge operator [7, 8, 9] incorporates contributions from neighboring elements, resulting in a sparse, non-diagonal matrix.
@@
The rest of the paper is organized as follows: II presents the related work. III provides preliminaries on aspects relevant to the proposed approach. IV presents the methodology, and V presents the experimental results. Finally, VI discusses the outcomes of this work and concludes this paper.
II Related Work
Recent advances in geometric deep learning have adapted convolutional neural networks from 2D to 3D geometry for classification and segmentation tasks. These methods adapt convolution operations for sampled 3D manifolds by exploiting relationships between mesh elements, i.e.ifnextchar.gobblevertices, edges and faces, to construct convolutional filters.
@@ II Related Work
III Preliminaries
III-A Triangular meshes & Discrete Operators
@@
All the above constructs can be naturally represented as matrices, as shown in table I. The matrix representations of the discrete exterior derivative operators $d_{0}\in\{-1,0,1\}^{n_{e}\times n_{v}}$ and $d_{1}\in\{-1,0,1\}^{n_{f}\times n_{e}}$ are simply the signed incidence matrices of the input meshes. The matrix representations of the Hodge Star operators, as mentioned above, depend on the discretization of the dual mesh.
@@
TABLE I: Mapping of different objects to matrices
@@ III-B Transformers
IV A Transformer Model with Hodge Attention
IV-A HodgeFormer Layer
@@ IV-B End-to-end Architecture
The proposed end-to-end architecture is configurable to combine operations on different mesh elements, i.e., vertices, edges, and faces, where each learned operator acts upon and updates specific k-forms, as depicted in table II. In the case that specific mesh elements are not selected to be updated, the corresponding Transformer layers are omitted. The size of the architecture in terms of trainable parameters depends on the number and configuration of used layers and the combination of mesh elements selected to operate upon.
@@
TABLE II: Overview of learnable operators $L_{v}$, $L_{e}$, $L_{f}$.
IV-C Sparse Attention
@@
IV-D Input Features and Positional Embeddings
@@
Specifically, we consider three types of input features: (i) coordinates, (ii) normals and (iii) areas: The vertex features $x_{v_{in}}$ consist of the point 3D coordinates, vertex normals, $n_{v_{i}}$ calculated as the weighted average of incident face normals $n_{f_{i}}$ and the vertex associated cell area, calculated as the weighted average of incident face areas. The edge features $x_{e_{in}}$ consist of the point 3D coordinates of edge vertices, and the vertices opposite to the respective edge, edge normals calculated as the average of incident vertex normals, and the lengths of the edges that belong to the incident faces. The face features $x_{f_{in}}$ consist of the point 3D coordinates of the face vertices ordered according to the face’s orientation, the face normal and the face area. These design choices aim to incorporate information associated both to the primal and the dual mesh. The effect of each type of input features is showcased in the ablation study in table IV.
IV-E Embedding Layer
The input features mentioned in section IV-D, $x_{v_{in}}$, $x_{e_{in}}$ and $x_{f_{in}}$ are transformed by an embedding layer into the corresponding matrices $x_{v}$, $x_{e}$ and $x_{f}$. The embedding layer (eq. 13), operates separately on each mesh element and involves two steps: (i) Feature aggregation over one-hop neighborhood defined by the mesh element’s adjacency structure and (ii) embedding to the target dimension via a feed-forward MLP.
@@ IV-E Embedding Layer
V Experimental Evaluation
V-A Evaluation Setup
@@
HodgeFormer layers use sparse attention, where each input element attends over $\sqrt{n}$ elements, consisting of local neighbors and random connections in an $4:1$ ratio. Plain Transformer Layers follow a standard implementation except for the incorporation of a linear attention mechanism as in  [32]. All MLP layers use ReLU non-linearities where required. We adopt the pre-LN architecture variant with residual connections, placing Layer Normalization (LayerNorm) before the input to each layer block. Following [28], we remove scale and bias terms from LayerNorm. Dropout is employed as a regularization mechanism both in the HodgeFormer and the Transformer layers. Finally, in all experiments we employ the neighbor-aware embedding layer as described in section IV-E.
V-B Experiments and Results
@@
The task of mesh classification is evaluated on the SHREC-11[33] and Cube Engraving [11] datasets. SHREC-11 is a dataset of 600 meshes representing 30 categories of 20 shapes each. Following prior work, we train on randomly selected splits of 10 samples per class and evaluate on the remaining data. Cube Engraving is a synthetic dataset of 2D shapes engraved on a randomly chosen face of a cube, comprising 4,381 shapes from 22 distinct categories. In both experiments, we optimize against a cross-entropy loss with a label smoothing factor of $0.2$. HodgeFormer achieves competitive results compared to other methods (table III). For SHREC-11, the best-performing model uses 6 HodgeFormer layers followed by 2 Transformer layers operating on all mesh elements. In contrast, for Cube Engraving, the top model also uses 6 HodgeFormer layers followed by 2 Transformer layers, but operates only on vertex features.
Figure 5: Mesh segmentation results on the Human Body test set.
@@
TABLE III: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[33] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of [34] with hard ground truth labels at faces, as well on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods.
@@
For the mesh segmentation task, our goal is to predict labels for every face of a mesh. We evaluate HodgeFormer architecture on four datasets: on the Human-part-segmentation dataset in its simplified version [34] with hard ground truth labels at faces, and the Vases, Chairs, and Aliens categories from the Shape COSEG dataset. In all experiments, we optimize against a class-weighted cross-entropy loss with an additional label smoothing factor of $0.2$. For each dataset, we report the highest performing variant of our architecture. For Human and Vases datasets, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, whereas for Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by 2 Transformer layers. As table III demonstrates, our approach performs comparatively to state of the art without spectral features, eigenvalue decomposition operations, or complex complementary structures.
V-C Model Variations Study
@@
To evaluate the importance of different architecture components and parameter choices, we varied our base model in different ways and measured its performance on a segmentation task. Specifically, in table V, we compare the main architecture layers HodgeFormer and a vanilla Transformer layer along with different embedding layers, i.e., the feed-forward and the neighbor embedding layers.
@@
In table IV, we test combinations of different input features based on their type by distinguishing between coordinates, normals, and areas. In Table  VI, we evaluate the effect of the number of neighbors taken into account by the sparse attention component on the performance of the HodgeFormer layers. For all experiments, we consider an architecture of 4 layers acting only on mesh vertices with embedding dimension 256 and hidden dimension 512, and a learning rate of 5e-4, and we report test accuracy results, averaged over 5 experiment trials with random training splits. We also consider $\sqrt{n}$ neighbors for each set of $n$ mesh elements with the exception of the experiment of table VI.
@@
In table V, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In table VI, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In table IV, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.
V-D Efficiency
@@
table VII compares the runtime performance of HodgeFormer with other recent methods on the human segmentation task, under typical training and testing configurations. HodgeFormer has an attractive runtime profile compared to other methods with fast training, high GPU utilization and no precomputation steps. It should be noted that a significant part of its execution time is spent in I/O and preprocessing, mostly in the BFS-based neighborhood extraction.
@@
TABLE IV: Comparing input feature effect on HodgeFormer architecture performance, evaluated on the COSEG Vases dataset.
@@
TABLE V: Ablation study on the effect of the Neighbor Embedding layer (section IV-E) evaluated on the COSEG Vases dataset. FNN refers to simple Feed-Forward Neural Network.
@@
TABLE VI: Ablation study on number of neighbors for the HodgeFormer Sparse Attention, evaluated on the COSEG Vases dataset.
@@
TABLE VII: Runtime performance comparison of HodgeFormer and other mesh-based methods on train and test sets of Human segmentation dataset, under typical training and testing configurations. We report the following metrics: (a) Execution time per epoch, (b) GPU time per epoch, (c) GPU time per batch, and (d) Peak GPU memory usage. All reported times are averaged over multiple epochs or batches, as appropriate.
VI Discussion, limitations and conclusion
@@ Appendix A End-to-end Architecture
A-A Multi-head Hodge Attention on vertices, edges and faces
This section presents a detailed analysis of the HodgeFormer attention mechanism, demonstrating how the multi-head Hodge Attention component operates across different mesh elements. By combining the Hodge star operator formulation (Eq. 12) with the attention-based operators (Eq. 6), the Hodge Laplacians $L_{v}$, $L_{e}$, and $L_{f}$ for vertices (0-forms), edges (1-forms), and faces (2-forms) are reformulated as the attention-based expressions given in equations 14, A-A, and 16, respectively.
@@ A-A Multi-head Hodge Attention on vertices, edges and faces
This formulation ensures that the attention mechanism respects the topological structure of the mesh while enabling information flow between different dimensional elements (vertices, edges, faces). The formation of these updated features is illustrated in Figures 8, 6, and 7, which visualize the computational flow described in equations (17), (18), and (19), respectively.
A-B Computational complexity and memory requirements
  1. 1.
  2. 2.
  3. 3.
  4. 4.
The resulting complexity is $O(nd^{2})+O(n^{1.5}d)$ with $d$ fixed and usually much smaller than $n$. In comparison, the eigendecomposition of a sparse $(n,n)$ matrix has complexity $O(kn^{2})$ for calculating the first $k$ eigenvectors. This holds for Laplacian positional embeddings as well as for spectral features such as Heat Kernel Signature (HKS). table VIII presents experiments evaluating the computational and memory requirements.
@@ A-B Computational complexity and memory requirements
[…]
@@
TABLE VIII: Metrics for different mesh sizes with respect to the number of vertices, for a 1-layer HodgeFormer (HF) end-to-end architecture (training and inference) as well as a standalone HodgeFormer layer (inference). The model operates on vertices with a latent embedding dimension $d=256$ and hidden MLP dimension of $d_{h}=512$. Measured on an Nvidia RTX 4090 GPU.
Appendix B Results
B-A Eigenfunctions of the corresponding learned operator
We compute the eigenfunctions of the corresponding learned operators and visualize them on the mesh geometry, providing valuable insights into how the HodgeFormer captures and processes geometric features at different scales and orientations. Figure 11 presents the first five eigenfunctions of a COSEG Aliens dataset sample.
@@ B-A Eigenfunctions of the corresponding learned operator
TABLE IX: Variation experiments comparing different strategies of mixing HodgeFormer with Transformer layers and their effect on the performance, evaluated on the COSEG Vases dataset.
B-B Ablation study on mixing strategies
This section examines the effect of mixing HodgeFormer and Transformer layers on segmentation performance. Table IX summarizes the performance on the COSEG Vases dataset under varying layer compositions. We denote $N_{H}$ as the number of HodgeFormer layers and $R_{H:T}=(N_{H}:N_{T})$ the ratio of the number of HodgeFormer layers to Vanilla Transformer layers. As shown in Table IX, increasing the number of Transformer layers while keeping the number of HodgeFormer layers fixed ($N_{H}=4$) shows a general improvement, with the best result (93.04%) obtained when mixing six layers in total ($R_{H:T}=4:2$).
B-C Learned attention maps
Figure 10 presents the learned attention maps for two sample meshes of COSEG Aliens dataset measured in layer 6 of the architecture.

### v2

@@ Abstract
1 Introduction
@@
In this work, we introduce a novel Transformer architecture inspired by the explicit construction of the Laplacian operator in Discrete Exterior Calculus (DEC) adjusted to triangular 3-dimensional meshes. This kind of connection is not novel by itself and was the main idea of HodgeNet by Smirnov and Solomon [33], where input features were used for learning Hodge Star operators as diagonal matrices. These matrices were then used for constructing Laplacian operators, which were then applied to input features in a spectral learning kind of approach. Specifically, we draw connections between discrete Hodge Star operators and Transformer-style attention mechanisms. While most common constructions of discrete Hodge Star operators are limited to diagonal matrices, these represent only some specific Hodge operator realizations. In fact, the Galerkin method, widely used in Finite Element Methods (FEM) literature [34], prescribes systematic ways to discretize various differential operators, including the Hodge operator. The key idea is to project onto carefully designed basis functions (e.g., piecewise linear or higher order elements) and account for the overlap of these basis functions. These overlaps are used to design the mass matrix, which acts as a discrete Hodge operator. Unlike the diagonal Hodge operator, which considers only self-contributions, the Galerkin Hodge operator [2, 27, 22] incorporates contributions from neighboring elements, resulting in a sparse, non-diagonal matrix.
@@
The rest of the paper is organized as follows: Sec. 2 presents the related work. Sec. 3 provides preliminaries on aspects relevant to the proposed approach. Sec. 4 presents the methodology, and 5 presents the experimental results. Finally, Sec. 6 discusses the outcomes of this work and concludes this paper.
2 Related Work
Recent advances in geometric deep learning have adapted convolutional neural networks from 2D to 3D geometry for classification and segmentation tasks. These methods adapt convolution operations for sampled 3D manifolds by exploiting relationships between mesh elements, i.e. vertices, edges and faces, to construct convolutional filters.
@@ 2 Related Work
3 Preliminaries
3.1 Triangular meshes & Discrete Operators
@@ 2. $i$
All the above constructs can be naturally represented as matrices, as shown in Tab. 1. The matrix representations of the discrete exterior derivative operators $d_{0}\in\{-1,0,1\}^{n_{e}\times n_{v}}$ and $d_{1}\in\{-1,0,1\}^{n_{f}\times n_{e}}$ are simply the signed incidence matrices of the input meshes. The matrix representations of the Hodge Star operators, as mentioned above, depend on the discretization of the dual mesh.
@@
Table 1: Mapping of different objects to matrices
@@ 3.2 Transformers
4 A Transformer Model with Hodge Attention
4.1 HodgeFormer Layer
@@ 4.2 End-to-end Architecture
The proposed end-to-end architecture is configurable to combine operations on different mesh elements, i.e., vertices, edges, and faces, where each learned operator acts upon and updates specific k-forms, as depicted in Tab. 2. In the case that specific mesh elements are not selected to be updated, the corresponding Transformer layers are omitted. The size of the architecture in terms of trainable parameters depends on the number and configuration of used layers and the combination of mesh elements selected to operate upon.
@@
Table 2: Overview of learnable operators $L_{v}$, $L_{e}$, $L_{f}$.
4.3 Sparse Attention
@@
4.4 Input Features and Positional Embeddings
@@
Specifically, we consider three types of input features: (i) coordinates, (ii) normals and (iii) areas: The vertex features $x_{v_{in}}$ consist of the point 3D coordinates, vertex normals, $n_{v_{i}}$ calculated as the weighted average of incident face normals $n_{f_{i}}$ and the vertex associated cell area, calculated as the weighted average of incident face areas. The edge features $x_{e_{in}}$ consist of the point 3D coordinates of edge vertices, and the vertices opposite to the respective edge, edge normals calculated as the average of incident vertex normals, and the lengths of the edges that belong to the incident faces. The face features $x_{f_{in}}$ consist of the point 3D coordinates of the face vertices ordered according to the face’s orientation, the face normal and the face area. These design choices aim to incorporate information associated both to the primal and the dual mesh. The effect of each type of input features is showcased in the ablation study in Tab. 4.
4.5 Embedding Layer
The input features mentioned in Sec. 4.4, $x_{v_{in}}$, $x_{e_{in}}$ and $x_{f_{in}}$ are transformed by an embedding layer into the corresponding matrices $x_{v}$, $x_{e}$ and $x_{f}$. The embedding layer (Eq. 13), operates separately on each mesh element and involves two steps: (i) Feature aggregation over one-hop neighborhood defined by the mesh element’s adjacency structure and (ii) embedding to the target dimension via a feed-forward MLP.
@@ 4.5 Embedding Layer
5 Experimental Evaluation
5.1 Evaluation Setup
@@
HodgeFormer layers use sparse attention, where each input element attends over $\sqrt{n}$ elements, consisting of local neighbors and random connections in an $4:1$ ratio. Plain Transformer Layers follow a standard implementation except for the incorporation of a linear attention mechanism as in  [13]. All MLP layers use ReLU non-linearities where required. We adopt the pre-LN architecture variant with residual connections, placing Layer Normalization (LayerNorm) before the input to each layer block. Following [30], we remove scale and bias terms from LayerNorm. Dropout is employed as a regularization mechanism both in the HodgeFormer and the Transformer layers. Finally, in all experiments we employ the neighbor-aware embedding layer as described in Sec. 4.5.
5.2 Experiments and Results
@@
The task of mesh classification is evaluated on the SHREC-11[20] and Cube Engraving [11] datasets. SHREC-11 is a dataset of 600 meshes representing 30 categories of 20 shapes each. Following prior work, we train on randomly selected splits of 10 samples per class and evaluate on the remaining data. Cube Engraving is a synthetic dataset of 2D shapes engraved on a randomly chosen face of a cube, comprising 4,381 shapes from 22 distinct categories. In both experiments, we optimize against a cross-entropy loss with a label smoothing factor of $0.2$. HodgeFormer achieves competitive results compared to other methods (Tab. 3). For SHREC-11, the best-performing model uses 6 HodgeFormer layers followed by 2 Transformer layers operating on all mesh elements. In contrast, for Cube Engraving, the top model also uses 6 HodgeFormer layers followed by 2 Transformer layers, but operates only on vertex features.
@@
Table 3: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[20] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of Milano et al. [26] with hard ground truth labels at faces, as well on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods.
@@
For the mesh segmentation task, our goal is to predict labels for every face of a mesh. We evaluate HodgeFormer architecture on four datasets: on the Human-part-segmentation dataset in its simplified version [26] with hard ground truth labels at faces, and the Vases, Chairs, and Aliens categories from the Shape COSEG dataset. In all experiments, we optimize against a class-weighted cross-entropy loss with an additional label smoothing factor of $0.2$. For each dataset, we report the highest performing variant of our architecture. For Human and Vases datasets, our highest performing model was a 66-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, whereas for Chairs and Aliens datasets, our highest performing model was a 66-layered HodgeFormer model followed by 2 Transformer layers. As Tab. 3 demonstrates, our approach performs comparatively to state of the art without spectral features, eigenvalue decomposition operations, or complex complementary structures.
Figure 5: Mesh segmentation results on the Human Body test set.
5.3 Model Variations Study
@@
To evaluate the importance of different architecture components and parameter choices, we varied our base model in different ways and measured its performance on a segmentation task. Specifically, in Tab. 5, we compare the main architecture layers HodgeFormer and a vanilla Transformer layer along with different embedding layers, i.e., the feed-forward and the neighbor embedding layers.
@@
In Tab. 4, we test combinations of different input features based on their type by distinguishing between coordinates, normals, and areas. In Table  6, we evaluate the effect of the number of neighbors taken into account by the sparse attention component on the performance of the HodgeFormer layers.
For all experiments, we consider an architecture of 4 layers acting only on mesh vertices with embedding dimension 256 and hidden dimension 512, and a learning rate of 5e-4, and we report test accuracy results, averaged over 5 experiment trials with random training splits. We also consider $\sqrt{n}$ neighbors for each set of $n$ mesh elements with the exception of the experiment of Tab. 6.
@@
In Tab. 5, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In Tab. 6, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In Tab. 4, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.
5.4 Efficiency
@@
Tab. 7 compares the runtime performance of HodgeFormer with other recent methods on the human segmentation task, under typical training and testing configurations. HodgeFormer has an attractive runtime profile compared to other methods with fast training, high GPU utilization and no precomputation steps. It should be noted that a significant part of its execution time is spent in I/O and preprocessing, mostly in the BFS-based neighborhood extraction.
@@
Table 4: Comparing input feature effect on HodgeFormer architecture performance, evaluated on the COSEG Vases dataset.
@@
Table 5: Ablation study on the effect of the Neighbor Embedding layer (Sec. 4.5) evaluated on the COSEG Vases dataset. FNN refers to simple Feed-Forward Neural Network.
@@
Table 6: Ablation study on number of neighbors for the HodgeFormer Sparse Attention, evaluated on the COSEG Vases dataset.
@@
Table 7: Runtime performance comparison of HodgeFormer and other mesh-based methods on train and test sets of Human segmentation dataset, under typical training and testing configurations. We report the following metrics: (a) Execution time per epoch, (b) GPU time per epoch, (c) GPU time per batch, and (d) Peak GPU memory usage. All reported times are averaged over multiple epochs or batches, as appropriate.
6 Discussion, limitations and conclusion
@@ Appendix A End-to-end Architecture
A.1 Multi-head Hodge Attention on vertices, edges and faces
This section presents a detailed analysis of the HodgeFormer attention mechanism, demonstrating how the multi-head Hodge Attention component operates across different mesh elements. By combining the Hodge star operator formulation (Eq. 12) with the attention-based operators (Eq. 6), the Hodge Laplacians $L_{v}$, $L_{e}$, and $L_{f}$ for vertices (0-forms), edges (1-forms), and faces (2-forms) are reformulated as the attention-based expressions given in equations 14, A.1, and 16, respectively.
@@ A.1 Multi-head Hodge Attention on vertices, edges and faces
This formulation ensures that the attention mechanism respects the topological structure of the mesh while enabling information flow between different dimensional elements (vertices, edges, faces). The formation of these updated features is illustrated in Figures 7, 8, and 6, which visualize the computational flow described in equations (17), (18), and (19), respectively.
@@
A.2 Computational complexity and memory requirements
  1. (a)
  2. (b)
  3. (c)
  4. (d)
The resulting complexity is $O(nd^{2})+O(n^{1.5}d)$ with $d$ fixed and usually much smaller than $n$. In comparison, the eigendecomposition of a sparse $(n,n)$ matrix has complexity $O(kn^{2})$ for calculating the first $k$ eigenvectors. This holds for Laplacian positional embeddings as well as for spectral features such as Heat Kernel Signature (HKS). Tab. 8 presents experiments evaluating the computational and memory requirements.
@@ A.2 Computational complexity and memory requirements
Table 8: Metrics for different mesh sizes with respect to the number of vertices, for a 1-layer HodgeFormer (HF) end-to-end architecture (training and inference) as well as a standalone HodgeFormer layer (inference). The model operates on vertices with a latent embedding dimension $d=256$ and hidden MLP dimension of $d_{h}=512$. Measured on an Nvidia RTX 4090 GPU.
@@
Table 9: Variation experiments comparing different strategies of mixing HodgeFormer with Transformer layers and their effect on the performance, evaluated on the COSEG Vases dataset.
[…]
Appendix B Results
B.1 Ablation study on mixing strategies
This section examines the effect of mixing HodgeFormer and Transformer layers on segmentation performance. Table 9 summarizes the performance on the COSEG Vases dataset under varying layer compositions. We denote $N_{H}$ as the number of HodgeFormer layers and $R_{H:T}=(N_{H}:N_{T})$ the ratio of the number of HodgeFormer layers to Vanilla Transformer layers. As shown in Table 9, increasing the number of Transformer layers while keeping the number of HodgeFormer layers fixed ($N_{H}=4$) shows a general improvement, with the best result (93.04%) obtained when mixing six layers in total ($R_{H:T}=4:2$).
B.2 Eigenfunctions of the corresponding learned operator
We compute the eigenfunctions of the corresponding learned operators and visualize them on the mesh geometry, providing valuable insights into how the HodgeFormer captures and processes geometric features at different scales and orientations. Figure 10 presents the first five eigenfunctions of a COSEG Aliens dataset sample. Figure 11 presents the learned attention maps for two sample meshes of COSEG Aliens dataset measured in layer 6 of the architecture.